# Reinforcement learning to learn quantum states for Heisenberg scaling accuracy

### v1

@@ II.2 RL agent
The RL agent consists of an actor and a critic, both implemented by feed-forward neural networks [see Fig. 1 (c)]. The both networks use three hidden layers. For each hidden layer, the critic has hundred nodes, and the actor has fifty nodes. We use ReLu for activation functions. The actor is to learn a policy of action. We construct the action space ${\cal A}$ by selecting $m$ values from each of the hyperparameter spaces ${\cal A}_{\sigma}=[0,\infty)$ and ${\cal A}_{\eta}=[0,\infty)$. So the number of actions the agent can choose from is $m^{2}=\left|\cal A\right|$. The output of actor is the policy $\pi_{\phi_{2}}=p(\sigma,\eta|o,\phi_{2})$, the probability distribution defined over the action space ${\cal A}$. The critic outputs an estimate of observation-value function $Q_{\phi_{1}}(o)$.
We train the agent to recommend the hyperparmaeters $\sigma$ and $\eta$ that enables the learning quantum states to be finished quickly. This is accomplished by giving a penalty of $-1$ for each time step, and no penalty for the time step when the learning quantum states is finished. In other words, the reward is given by
@@
For the training, we use the Actor-Critic algorithm, and collect training datasets using a tree search guided by the policy. See Appendix C for detailed description.
Table 1: The settings used to train the RL agent. The space of action is given by ${\cal A}={\cal A}^{4}_{\sigma}\times{\cal A}^{4}_{\eta}$. The action repetition time $t_{\text{rep}}$ is determined by the parameters $(t_{l},t_{u},T_{\text{th}})$. For $4$- and $5$-qubit, we perform the learning random pure states with the RL agent trained using the $3$-qubit states.
@@ III.1 Training RL agent
[…]
We present the setting of HEA, the number of samples, and the space of action in Table 1. While the number of parameters of unitary operator required to prepare an $N$-qubit state is $4^{N}$, we show that the HEA can learn the random quantum states with fewer parameters: For the learning $2$- and $3$-qubit states, we use the HEAs which have $12$ and $54$ parameters, respectively. We will also show that HEA of $10$ layers can learn random $4$- and $5$-qubit states. Moreover, to estimate the gradient of HEAs, we use fewer samples than the parameters of HEA circuits. For $2$- and $3$-qubit states, the number of samples is set to $10$ and $30$, respectively.
Figure 3: (a) The scheme of action repetition strategy (ARS). If the ARS is not applied, the agent $\pi$ draws a new action $a_{i}$ based on the observation $o_{t}$ at each time step. With the ARS, the agent repeats an action chosen during $t_{\text{rep}}$ time steps. The example shows the case of $t_{\text{rep}}=3$. (b) Learning random $1$-qubit states with and without ARS. The red solid line is obtained by using the ARS with setting $(t_{u},t_{l},T_{\text{th}})=(50,1,100)$. The gray solid line is obtained without the ARS. The dashed line is the baseline value (ES only).
The results of training RL agent are illustrated in Fig. 2 (a). For these results, we set the target success count to ${\cal C}_{\text{target}}=10^{4}$, and randomly initialize the parameters of HEAs. We compare these results to the performance obtained without the RL, i.e., using only a single action for the entire process of learning quantum states. We regard results obtained from an action which gives the lowest total success count as baseline values, and train the RL agent until it outperforms the baseline values. The baselines are selected based on data in Appendix. D.
@@
The number of actions in the extended space is $169$. Except for the number of actions, we use the same settings for the $2$-qubit case listed in Table. 1. The result is shown in the rightmost column of Fig. 2 (a). By extending the action space, we can attain the lower values of average total success count and infidelity than the case of using the $16$ actions. The average total success count is $3.593\times 10^{5}$ and the average fidelity is $0.9998$. To achieve these results, on average, $433$ time steps are required. While the overall performance of learning $2$-qubit states is improved by extending the action space, the agent requires more episodes to be trained.
Fig. 2 (b) shows the policy of trained actor. These policies imply that the trained RL agent takes a strategy; if the success count is small, the agent recommends the large values for $\sigma$ and $\eta$ to enable the ES to search in a wide range of parameter space; if the success count is large, the agent reduces the values of recommended hyperparameter to make the learning quantum states to converge.
Figure 4: Learning random pure states with trained RL agent. We consider the target success counts from $10^{1}$ to $10^{4}$. The dots in (a)–(c) and (e) represent the average total success count $\langle{\cal C}_{\text{total}}\rangle$ and the average infidelity $\langle\bar{f}\rangle$ for the learning $1$-qubit (red), $2$-qubit (green), and $3$-qubit (blue) states. We denote the results obtained using the trained RL agent, referred to as ES$+$RL. The black triangles are the average values obtained without the RL agent, which are denoted by ES. For the results of ES, we use an action $a_{\text{base}}$ which gives the baseline values. SSML denotes the results obtained by the single-shot measurement learning. (d) $\Delta\langle{\cal C}_{\text{total}}\rangle=\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES}}}-\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES+RL}}}$ is the amount of average total success counts reduced by using the trained RL agent. (e) The purple and black dots are results of generalizing the RL agent trained using $3$-qubit states to learning random $4$- and $5$-qubit states, respectively. The dashed lines represent the scaling of statistical limit, $\bar{f}\sim{\cal O}({\cal C}^{-1}_{\text{total}})$.
@@ III.2 Action repetition strategy
To obtain these results, we use a method, called action repetition strategy (ARS), that repeats an action given by the RL agent for a certain period of time steps $t_{\text{rep}}$. The schematic of ARS is illustrated in Fig. 3 (a). For each RL episode $T$, $t_{\text{rep}}$ is given by
@@
The ARS is a method inspired by curriculum learning, a technique that trains a machine learning model by gradually increasing the difficulty of problems [44]. Note that the curriculum learning belongs to the category of transfer learning, which reuses a model or knowledge obtained from a task to learn a related task [70, 71, 72]. One of factors that determines the difficulty of RL problem is the depth of decision process. In general, the deeper the decision process, the more difficult the problem is considered to be. The depth of decision process of our problem $t_{H}$ in the early episodes is about $10^{3}$, $10^{4}$, and $1.5\times 10^{4}$ for the learning $1$-, $2$-, and $3$-qubit states, respectively. With the ARS, we can reduce the depth of decision process to $\left\lfloor t_{H}/t_{\text{rep}}\right\rfloor$, which is shallower than the original depth. Specifically, according to the settings in Table 1, the decision process can be reduced to tens of steps in the beginning of the RL training. This leads to the significant decrease of difficulty of problem. As the RL training proceeds, the difficulty of problem increases again with the decrease of $t_{\text{rep}}$. Thus, by controlling the action repetition time, the ARS enables the RL agent to experience problems of various difficulty levels during the training process.
For the random $1$-qubit states, we compare the performance obtained with and without the ARS in Fig 3 (b). With the ARS, the RL agents can converge faster than the case without the ARS. For the random $2$- and $3$-qubit states, the training RL agent without the ARS requires computational resources far exceed our capacity. In some intervals of RL episodes, the difficulty of our problem does not gradually increase since the halting time step $t_{H}$ also decreases together with the action repetition time. More sophisticated curriculum for the repetition time may improve the performance of meta-learning, but it is beyond the scope of this work and remains an open question.
Figure 5: Learning an entangled $5$-qubit pure state $\hat{\varrho}_{SC}$ with the RL agent trained using the random $3$-qubit states. (a) Re$\left(\hat{\varrho}_{SC}\right)$ represents the real parts of density matrix of $5$-qubit state. The imaginary parts of density matrix are zero. (b) The reconstructed state $\hat{\varrho}_{\text{est}}$ of the $5$-qubit state. Re$\left(\hat{\varrho}_{\text{est}}\right)$ and Im$\left(\hat{\varrho}_{\text{est}}\right)$ represent real and imaginary parts of density matrix $\hat{\varrho}_{\text{est}}$, respectively. The reconstructed state achieves the fidelity $f\approx 0.9989$. (c) The changes of infidelity during the entire learning process.
@@ III.3 The scaling of infidelity
Using the RL agents trained with the target success count ${\cal C}_{\text{target}}=10^{4}$, we perform the learning a hundred random pure states by varying the target success counts from $10^{1}$ to $10^{4}$. The results are shown in Fig. 4 (a)–(c). We compare these results to the baseline values obtained by an action $a_{\text{base}}$ (see Appendix D). Fig. 4 (d) shows the amount of total success counts saved by the RL agent according to the increase of the number of qubits $N$ and target success count ${\cal C}_{\text{target}}$. The amount of total success counts saved is represented by $\Delta\langle{\cal C}_{\text{total}}\rangle:=\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES}}}-\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES+RL}}}$, where $\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES}}}$ is obtained using $a_{\text{base}}$ without the RL agent, and $\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES+RL}}}$ is the average total success count obtained with the trained RL agent. In the target success count $10^{4}$, $\Delta\langle{\cal C}_{\text{total}}\rangle$ is given by $3.274\times 10^{4}$, $1.995\times 10^{5}$, and $2.818\times 10^{6}$ for $1$-, $2$-, and $3$-qubit states, respectively. This leads to enhancement in the scaling of average infidelity $\bar{f}$ versus the total success counts ${\cal C}_{\text{total}}$. The scaling nearly achieves the Heisenberg limit [45], $\bar{f}\sim{\cal O}({\cal C}^{-1}_{\text{total}})$, which the standard QST method cannot attain without additional information about quantum states [73].
We curve fit the results of Fig. 4 (a)–(c) with $\bar{f}=\alpha{\cal C}_{\text{total}}^{-\beta}$. The scaling factor is given by $\beta\approx 0.948$, $1.161$, and $1.184$ for the learning $1$-, $2$-, and $3$-qubit states, respectively. Our method shows the similar results to previous works on learning $1$-qubit states [74, 75, 76, 25, 26], where the scaling factors lie in the range $\beta\in[0.9,1.0)$. Typically, the SSML shows the scaling of $\beta\approx 0.972$ in our simulation, which is consistent with the results in their works [25, 26]. (In the multi-qubit cases, we fail to find hyeperameter settings of SSML to achieve the optimal scaling.) Our result obtained with the RL agent achieves the similar scaling factor while using fewer total success count than the SSML as shown in Fig. 4 (a).
@@ III.4 Generalization of RL agent
We apply the RL agent trained with the random $3$-qubit states to learning a hundred random $4$- and $5$-qubit states. The results are illustrated in Fig. 4 (e). For the target success count ${\cal C}_{\text{target}}=10^{4}$, the average of total success count and infidelity are given by $\langle{\cal C}_{\text{total}}\rangle\approx 7.261\times 10^{6}$ and $\bar{f}\approx 6.561\times 10^{-4}$ for $4$-qubits, and $\langle{\cal C}_{\text{total}}\rangle\approx 1.291\times 10^{8}$ and $\bar{f}\approx 9.311\times 10^{-4}$ for $5$-qubits. The scaling factor of infidelity achieves $\beta=-1.189$ and $\beta=-0.829$ for $4$- and $5$-qubit, respectively. For the multi-qubit cases including $2$- and $3$-qubit results, our method gives the similar results to the self-guided quantum state tomography (SGQT) which shows the scaling factors in the range $\beta\in(0.80,1.05)$ up to $10$-qubits [76].
@@
By using the trained RL agent, we learn a specific $5$-qubit state, and reconstruct the learned state. As illustrated in Fig. 5 (a), the density matrix of $5$-qubit state $\hat{\varrho}_{SC}$ is given by a Shen-Castan matrix of which element is determined by a real-valued function $\exp[-(\left|m-16.5\right|+\left|n-16.5\right|+1)/10]/K$ for $m,n=1,\ldots,32$ and $K=0.00366$. (A Shen-Castan matrix is a filer used for edge detection in image processing [77].) This quantum state is pure, i.e., $\mbox{Tr}\hat{\varrho}_{SC}^{2}=1$, and the five qubits are entangled. For a reduced state of $i$-th subsystem $\hat{\varrho}_{i}$, Von Neumann entropy ${S}_{i}:=-\mbox{Tr}\hat{\varrho}_{i}\log_{2}\hat{\varrho}_{i}$ has nonzero value: ${S}_{1}=0.640$, ${S}_{2}=0.547$, ${S}_{3}=0.230$, ${S}_{4}=0.080$, and ${S}_{5}=0.025$. In the learning of the $5$-qubit state, we achieve the fidelity ${f}\approx 0.9989$ by using $6.62\times 10^{7}$ success count. The result of state reconstruction is illustrated in Fig. 5 (b). Fig. 5 (c) shows the changes of infidelity during the learning process. The policy of RL agent induces exploration in the early time steps, and, as the ES approaches the optimal point, the agent focuses on exploitation to improve the quality of solution.

### v2

@@ II.2 RL agent
The RL agent consists of an actor and a critic, implemented by the feed-forward neural networks [see Fig. 1 (c)]. The both networks use three hidden layers. For each hidden layer, the critic has hundred nodes, and the actor has fifty nodes. We use ReLu for activation functions. The actor is to learn a policy of action. We construct the action space ${\cal A}$ by selecting $m$ values from each of the hyperparameter spaces ${\cal A}_{\sigma}=[0,\infty)$ and ${\cal A}_{\eta}=[0,\infty)$. So the number of actions the agent can choose from is $m^{2}=\left|\cal A\right|$. The output of actor is the policy $\pi_{\phi_{2}}=p(\sigma,\eta|o,\phi_{2})$, the probability distribution defined over the action space ${\cal A}$. The critic outputs an estimate of observation-value function $Q_{\phi_{1}}(o)$.
We train the agent to recommend the hyperparmaeters $\sigma$ and $\eta$ that enables the learning of quantum states to be finished quickly. This is accomplished by giving a penalty of $-1$ to the RL agent at each time step, and no penalty for the final time step. Thus, the reward is given by
@@
For the training, we use the Actor-Critic algorithm, and collect training datasets using a tree search guided by the policy. See Appendix C for details.
Table 1: Settings used to train the RL agent. The space of action is given by ${\cal A}={\cal A}^{4}_{\sigma}\times{\cal A}^{4}_{\eta}$. The action repetition time $t_{\text{rep}}$ is determined by the parameters $(t_{l},t_{u},T_{\text{th}})$. For $4$- and $5$-qubit, we perform learning of random pure states with the RL agent trained using the $3$-qubit states.
@@ III.1 Training RL agent
We apply our meta-learning scheme to the learning of random pure states. The random states are prepared by applying unitary operators $\hat{V}$ drawn from Haar measure to the initial states $\left|0\right>^{\otimes N}$ [69]. To measure these states, we perform measurements in the computational basis, where $|0\rangle\langle 0|^{\otimes N}$ corresponds to the basis of the success outcome $|s\rangle\langle s|$. For each RL episode $T$, we prepare a new set of a thousand problem instances, and perform the learning of quantum states to collect datasets for RL. We obtain the following results through simulations implemented by PyTorch [70], running on four NVIDIA Tesla A100100100100 graphic processing units (GPUs). The code is available at Ref. [71].
We present the setting of the HEAs, the number of samples, and the space of action in Table 1. While the number of parameters of a unitary operator required to prepare an $N$-qubit state is $4^{N}$, we show that the HEA can learn the random quantum states with fewer parameters: For the learning $2$- and $3$-qubit states, we use the HEAs which have $12$ and $54$ parameters, respectively. We will also show that HEA of $10$ layers can learn random $4$- and $5$-qubit states. Moreover, to estimate the gradient of HEAs, we use fewer samples than the parameters of HEA circuits. For $2$- and $3$-qubit states, the number of samples is set to $10$ and $30$, respectively. These results show the expressivity of the HEAs.
Figure 3: (a) Scheme of action repetition strategy (ARS). Without the ARS ($t_{\text{rep}}=0$), the agent $\pi$ draws a new action $a_{i}$ at each time step based on the observation $o_{t}$. With the ARS ($t_{\text{rep}}\neq 0$), the agent repeats a selected action for $t_{\text{rep}}$ consecutive time steps. The illustration depicts the cases where $t_{\text{rep}}=0$ and $t_{\text{rep}}=3$. (b) Learning random $1$-qubit states with and without the ARS. The red solid line represents the result obtained using the ARS with setting $(t_{u},t_{l},T_{\text{th}})=(50,1,100)$. The gray solid line corresponds to the case without the ARS. The dashed line indicate the baseline performance (ES only).
The results of training RL agent are illustrated in Fig. 2 (a). For these results, we set the target success count to ${\cal C}_{\text{target}}=10^{4}$, and randomly initialize the HEAs. We compare these results to the baseline values, the performance obtained without the RL, i.e., using only a single action for the entire process of learning quantum states. We regard the results obtained from an action which gives the lowest total success count as the baseline values, and train the RL agent until it outperforms the baseline values. The baselines are selected from the data in Appendix D.
@@
The number of actions in the extended space is $169$. Except for the number of actions, we use the same settings for the $2$-qubit case listed in Table. 1. The result is shown in the rightmost column of Fig. 2 (a). By extending the action space, we can attain the lower values of average total success count and infidelity than the case of using the $16$ actions; the average total success count is $3.593\times 10^{5}$ and the average fidelity is $0.9998$. To achieve these results, on average, $433$ time steps are required. While the overall performance of learning $2$-qubit states is improved by extending the action space, the agent requires more episodes to be trained.
Fig. 2 (b) shows the policy of trained actor. These policies imply that the trained RL agents take a strategy; if the success count is small, the agent recommends the large values for $\sigma$ and $\eta$ to enable the ES to search in a wide range of parameter space; if the success count is large, the agent reduces the values of the recommended hyperparameters to make the learning of quantum states converge.
Figure 4: Learning random pure states with the trained RL agent. We consider the target success counts from $10^{1}$ to $10^{4}$. The dots in (a)–(c) and (e) represent the average total success count $\langle{\cal C}_{\text{total}}\rangle$ and the average infidelity $\langle\bar{f}\rangle$ for the learning $1$-qubit (red), $2$-qubit (green), and $3$-qubit (blue) states. We denote the results obtained using the trained RL agent, labeled as ES$+$RL. The black triangles, labeled as ES, are the average values obtained without the RL agent. For the results of ES, we use an action $a_{\text{base}}$ which gives the baseline values. SSML represents the results obtained by the single-shot measurement learning. (d) $\Delta\langle{\cal C}_{\text{total}}\rangle=\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES}}}-\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES+RL}}}$ is the amount of average total success counts reduced by using the trained RL agent. (e) The purple and black dots are results of generalizing the RL agent trained using $3$-qubit states to learning random $4$- and $5$-qubit states, respectively. The dashed lines represent the scaling of Heisenberg limit, $\bar{f}\sim{\cal O}({\cal C}^{-1}_{\text{total}})$.
@@ III.2 Action repetition strategy
To obtain these results, we use a method, called action repetition strategy (ARS), that repeats an action selected by the RL agent for consecutive time steps $t_{\text{rep}}$. The schematic of ARS is illustrated in Fig. 3 (a). For each RL episode $T$, $t_{\text{rep}}$ is given by
@@
The ARS is a method inspired by curriculum learning, a technique that trains a machine learning model by gradually increasing the difficulty of problems [44, 45, 46]. Note that the curriculum learning belongs to the category of transfer learning, which reuses a model or knowledge obtained from a task to learn a related task [72, 73, 74]. One of factors that determines the difficulty of RL problem is the depth of decision process. In general, the deeper the decision process, the more difficult the problem is considered to be. The depths $t_{H}$ of decision processes of our problems in the early episodes are about $10^{3}$, $10^{4}$, and $1.5\times 10^{4}$ for the learning of $1$-, $2$-, and $3$-qubit states, respectively. With the ARS, we can reduce the depths of decision processes to $\left\lfloor t_{H}/t_{\text{rep}}\right\rfloor$, which are shallower than the original depths. Specifically, according to the settings in Table 1, the decision processes can be reduced to tens of steps in the beginnings of the RL trainings. These lead to the significant decrease of difficulties of the problems. As the RL training proceeds, the difficulties of the problems increase again with the decrease of $t_{\text{rep}}$. Thus, by controlling the action repetition time, the ARS enables the RL agent to experience problems of various difficulty levels during the training process.
For the random $1$-qubit states, we compare the performance obtained with and without the ARS in Fig 3 (b). With the ARS, the RL agents can converge faster than the case without the ARS. For the random $2$- and $3$-qubit states, the training of RL agent without the ARS requires computational resources far exceed our capacity. In some intervals of RL episodes, the difficulties of our problems do not gradually increase since the halting time steps $t_{H}$ also decrease together with the action repetition times. More sophisticated curriculum for the repetition time may improve the performance of meta-learning, but it is beyond the scope of this work and remains an open question.
[…]
@@ III.3 The scaling of infidelity
Using the RL agents trained with the target success count ${\cal C}_{\text{target}}=10^{4}$, we perform learning of a hundred random pure states by varying the target success counts from $10^{1}$ to $10^{4}$. The results are shown in Fig. 4 (a)–(c). We compare these results to the baseline values obtained by an action $a_{\text{base}}$ (see Appendix D). Fig. 4 (d) shows the amount of total success counts saved by the RL agent according to the increase of the number of qubits $N$ and target success count ${\cal C}_{\text{target}}$. $\Delta\langle{\cal C}_{\text{total}}\rangle:=\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES}}}-\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES+RL}}}$ quantifies the amount of total success counts saved by the RL agent, where $\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES}}}$ is obtained using $a_{\text{base}}$ without the RL agent, and $\langle{\cal C}_{\text{total}}\rangle_{\text{{\bf ES+RL}}}$ is the average total success count obtained with the trained RL agent. In the target success count $10^{4}$, $\Delta\langle{\cal C}_{\text{total}}\rangle$ is given by $3.274\times 10^{4}$, $1.995\times 10^{5}$, and $2.818\times 10^{6}$ for $1$-, $2$-, and $3$-qubit states, respectively. This leads to enhancement in the scaling of average infidelity $\bar{f}$ with respect to the total success count ${\cal C}_{\text{total}}$.
The infidelity of our method scales according to the Heisenberg limit [47], $\bar{f}\sim{\cal O}({\cal C}^{-1}_{\text{total}})$. We curve fit the results of Fig. 4 (a)–(c) with $\bar{f}=\alpha{\cal C}_{\text{total}}^{-\beta}$. The scaling factor is given by $\beta\approx 0.948$, $1.161$, and $1.184$ for the learning $1$-, $2$-, and $3$-qubit states, respectively. Our method shows the similar results to previous works on the learning of $1$-qubit states [76, 77, 78, 25, 26], where the scaling factors lie in the range $\beta\in[0.9,1.0)$. Typically, the SSML shows the scaling of $\beta\approx 0.972$ in our simulation, which is consistent with the results in their works [25, 26]. (In the multi-qubit cases, we fail to find the settings of SSML to achieve the optimal scaling.) Our result obtained with the RL agent achieves the similar scaling factor while using fewer total success count than the SSML as shown in Fig. 4 (a). The standard QST method cannot achieve Heisenberg-limit scaling without additional information about quantum states, with its scaling limit given by $\beta=0.75$ [75].
@@ III.4 Generalization of RL agent
We apply the RL agent trained with the random $3$-qubit states to learning of a hundred random $4$- and $5$-qubit states. The results are illustrated in Fig. 4 (e). For the target success count ${\cal C}_{\text{target}}=10^{4}$, the average of total success count and infidelity are given by $\langle{\cal C}_{\text{total}}\rangle\approx 7.261\times 10^{6}$ and $\bar{f}\approx 6.561\times 10^{-4}$ for $4$-qubits, and $\langle{\cal C}_{\text{total}}\rangle\approx 1.291\times 10^{8}$ and $\bar{f}\approx 9.311\times 10^{-4}$ for $5$-qubits. The scaling factor of infidelity achieves $\beta=-1.189$ and $\beta=-0.829$ for $4$- and $5$-qubit, respectively. For the multi-qubit cases including the $2$- and $3$-qubit results, our method gives the similar results to the self-guided quantum state tomography (SGQT) which shows the scaling factors in the range $\beta\in(0.80,1.05)$ up to $10$-qubits [78]. In Appendix E, we investigate the performance of our scheme for mixed states.
@@
The trained RL agent can generalize to learning of a specific $5$-qubit state. As illustrated in Fig. 5 (a), the density matrix of the $5$-qubit state $\hat{\varrho}_{SC}$ is given by a Shen-Castan matrix, where each element is determined by a real-valued function $\exp[-(\left|m-16.5\right|+\left|n-16.5\right|+1)/10]/K$ for $m,n=1,\ldots,32$ and $K=0.00366$. (A Shen-Castan matrix is a filer used for edge detection in image processing [79].) This quantum state is pure, i.e., $\mbox{Tr}\hat{\varrho}_{SC}^{2}=1$, and the five qubits are entangled. For a reduced state of $i$-th subsystem $\hat{\varrho}_{i}$, Von Neumann entropy ${S}_{i}=-\mbox{Tr}\hat{\varrho}_{i}\log_{2}\hat{\varrho}_{i}$ has nonzero value: ${S}_{1}=0.640$, ${S}_{2}=0.547$, ${S}_{3}=0.230$, ${S}_{4}=0.080$, and ${S}_{5}=0.025$. (Note that subsystems of quantum states sampled from Haar measure exhibit nearly maximal entropy [80].) In the learning of the $5$-qubit state, we achieve fidelity ${f}\approx 0.9989$ by using $6.62\times 10^{7}$ success count. The result of state reconstruction is illustrated in Fig. 5 (b). Fig. 5 (c) shows the changes of infidelity throughout the learning process. The policy of the RL agent induces exploration in the early time steps, and, as the ES approaches the optimal point, the agent focuses on exploitation to improve the quality of solution.
Figure 6: Comparison between our meta-learning method and the quantum state tomography (QST). The results of learning a hundred random pure states. The infidelity and the number of shots are represented on a logarithmic scale in base $10$. The red, green, blue, purple, and black dots represent data obtained using our meta-learning. These are equivalent to the data labeled as ES+RL in Fig. 4. The square dots indicate results from SSML. The solid, dashed, and dot-dashed lines depict the results of QST based on the maximum likelihood estimation. For estimating random $N$-qubit states, the QST considers the $14^{N}-$ 1 measurement settings, where each setting uses $n$ copies of the states (or $n$ shots).
III.5 Comparison with the quantum state tomography
We compare our results with the performance of standard method, quantum state tomography (QST), based on maximum likelihood estimation [4, 5]. Fig. 6 shows the results of QST, demonstrating that the RL-driven meta-learning scheme learns the random quantum states more efficiently in terms of resource usage than the standard method.
To estimate a hundred random $N$-qubit states with QST, we assume a parametric model $\hat{\varrho}(\vec{r})=\vec{r}\cdot{\cal P}_{N}/2^{N}$, where $\vec{r}$ is a generalized Bloch vector of $4^{N}$ dimension, and ${\cal P}_{N}$ is a string of Pauli operators, $\{I,\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z}\}^{\otimes N}$. The first element of parameter vector is fixed to one, and the $14^{N}-$ 1 elements of the initial parameter vector $\vec{r}_{0}$ are randomly chosen. We consider $14^{N}-$ 1 measurements of which outcomes are associated with projectors given by $\{M_{I},M_{\pm X},M_{\pm Y},M_{\pm Z}\}^{\otimes N}/M^{\otimes N}_{I}$, where $M_{I}=I/2$, $M_{\pm X}=(I\pm\hat{\sigma}_{x})/2$, $M_{\pm Y}=(I\pm\hat{\sigma}_{y})/2$, and $M_{\pm Z}=(I\pm\hat{\sigma}_{z})/2$. Each measurement uses $n$ copies of unknown states (or equivalently $n$ shots). See Appendix F for a detailed description of the QST method.